OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = –3 and a5 = –768?

OpenStudy (anonymous):

i know how to do the equation i just need someone to ell me what step im doing wrong

OpenStudy (anonymous):

first if first term = a then 5th term = ar^4 so ar^4 / a = r^4 = -768 / -3 = 256 so r = 4 seventh term = ar^6 = -3 * 4^6 = -3* 4096 = - 12288